Find an question in the form y=ax^2+bx+c for the parabola passing through the points. (-4,-123), (3,-102), (-1,-6)
Answer by TimothyLamb(4379)

y = axx + bx + c
---
(-4,-123):
16a - 4b + c = -123
---
(3,-102):
9a + 3b + c = -102
---
(-1,-6):
a - b + c = -6
---
16a - 4b + c = -123
9a + 3b + c = -102
a - b + c = -6
---
put the system of linear equations into standard form
---
substitute:
x for a
y for b
z for c
---
16x - 4y + z = -123
9x + 3y + z = -102
x - y + z = -6
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = -9
y = -6
z = -3
---
substitute:
a for x
b for y
c for z
---
solution:
a = -9
b = -6
c = -3
---
answer:
y = axx + bx + c
y = -9xx - 6x - 3
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---